# Robustness-aware Automatic Prompt Optimization

### v2

@@ 6 Conclusion
In this paper, we introduce BATprompt, a novel approach designed to optimize prompts’ performance on perturbed datasets by combining prompt optimization algorithms with adversarial training. By leveraging LLMs to simulate gradients, BATprompt enables adversarial attacks and iterative optimization. Our extensive experiments demonstrate the superiority of BATprompt across multiple tasks and various types of perturbations. Compared to existing prompt optimization methods, BATprompt shows significant improvements in handling fragile perturbations and good generalization.
@@ 7 Limitation
We employed a training approach inspired by FGSM, which, while effective, could be enhanced by more advanced or specialized methods suited to particular perturbation types or task requirements. Additionally, although our work covers a range of widely used tasks and datasets, future exploration may involve expanding into more specialized or innovative application domains. Our reliance on black-box model access poses another constraint, as deeper access to model-internal gradients could potentially offer richer interpretability and more granular refinements during adversarial training. Pursuing open-source architectures or collaboration with model providers may therefore create new opportunities to improve both robustness and understanding of how perturbations interact with large language models.
Appendix A Weakness of Tasks
We used the same dataset as in section 5, and GPT-3.5-turbo was used as the test model. When constructing the perturbed data set, we also adopt the data generation way in section 5.
A.1 Language Understanding
The  Figure 5 illustrates the performance scores for a text classification task under various types of perturbations. Six datasets were analyzed, with each dataset employing two system prompts: manual instruction and natural instruction to guide the tasks. The scores for each task were averaged to determine the specific impact of each type of perturbation under different prompts. The results reveal that under manual instruction, the S1 and C3 perturbations have minimal impact on the final task scores. Consequently, these two types of perturbations were excluded from further consideration during selection.
A.2 Language Generation
The detailed results for text summarization and text simplification tasks are presented in the  Figure 6 and  Figure 7. Using manual instruction, we computed the Rouge-1/2/L scores and SARI values for the two datasets, XSum and ASSET, under various perturbations. For the text summarization task, the results indicate that the perturbations C3 and W2 do not significantly affect the Rouge-2 and Rouge-L scores. Therefore, these perturbations were not identified as weaknesses in the text summarization dataset. Similarly, for the text simplification task, perturbations C1, C2, and C3 have no negative impact on the SARI values, and thus, they were not considered weaknesses for the text simplification dataset.
Figure 5: An example of adding a perturbation to an input, where the top half indicates that the prompt gets the correct output under normal input, and the bottom half indicates that the prompt gets the wrong result under perturbed input
Figure 6: An example of adding a perturbation to an input, where the top half indicates that the prompt gets the correct output under normal input, and the bottom half indicates that the prompt gets the wrong result under perturbed input
Figure 7: An example of adding a perturbation to an input, where the top half indicates that the prompt gets the correct output under normal input, and the bottom half indicates that the prompt gets the wrong result under perturbed input
Appendix B Adversarial Attack Results
In our adversarial attack process, we use both semantic and structural similarity of the text before and after the attack as constraints. This ensures that, while the adversarial attack negatively impacts the target task, the resulting text maintains a high degree of similarity to the original. Specifically, for P1-type attacks, we calculate Levenshtein distance, while for P2-type attacks, we compute semantic similarity. The  Table 6 the similarity between the attacked text and the original text.
@@ Appendix B Adversarial Attack Results
Table 6: The averagesemantic similarity and Levenshtein distance before and after attack in 3 tasks, 8 datasets
From the table, we can observe that for long text inputs, such as those in the XSum dataset, the Levenshtein distance and semantic similarity between the original and perturbed text remain above 98% for both types of perturbations (P1 and P2). In contrast, for shorter text inputs, such as those in the Asset dataset and the six language understanding tasks, even small changes can significantly impact the similarity measures. However, our experiments show that the Levenshtein distance similarity for P1 perturbed data remains above 90%, while the semantic similarity for P2 perturbed data stays above 80%, except for the SST-5. This indicates that our attack effectively preserves the essential information of the original sentence while introducing controlled perturbations.
Appendix C Experiment Settings
C.1 Datasets
The  Table 7 shows the statistics of datasets we made for language understanding, text simplification, and text summarization tasks under different perturbations. Each dataset contains multiple sub-datasets, each of which is a specific class of perturbation for which the dataset feels weak. For instance, Xsum dataset contains 7 sub-datasets including Xsum under C2, C1, S1, W3, S3, S2, W1 perturbations.
Dataset Task |Train| |Test|
SST-2 language understanding {positive, negative} 1000 200
CR language understanding {positive, negative} 1000 200
MR language understanding {positive, negative} 1000 200
SST-5 language understanding {terrible, bad, okay, good, great} 1000 200
AG’s News language understanding {World, Sports, Business, Tech} 1000 200
TREC language understanding {Description, Entity, Expression, Human, Location, Number} 1000 200
ASSET Text Simplification 787 200
XSUM Text Summarization 520 200
Table 7: Statistics of our generating datasets for language understanding task and language generation task used in this work.
@@ C.1 Datasets
C.2 Hpyper Parameters
Our BATprompt algorithm is based on GPT-3.5-turbo for generation, with a total of 5 adversarial training iterations. During the adversarial attack phase, we set the number of iterative attacks to 3. In the combined adversarial attack and prompt optimization stages, we configured GPT-3.5-turbo with a Top-p value of 0.95 and a temperature of 1 to ensure both the robustness of the adversarial attack and the diversity of the generated outputs. For the testing phase, we set Top-p to 1 and temperature to 0, ensuring that the model produces consistent, fixed outputs.
@@ C.2 Hpyper Parameters
C.3 Template
In this section, we give a complete template for the perturbation adding phase.  Figure 8, the prompt optimization phase.  Figure 9 provides a detailed illustration of the gradients generated during the optimization phase.  Figure 10 offers a comprehensive explanation of the prompt generation process, where gradients guide the LLM, and highlights how prompt richness is enhanced through the rewriting process.
Here is the Implementation details of BATprompt.
@@ C.3 Template
Figure 11: BATprompt is based on the details of the LLM implementation, with the top half representing the adversarial attack phase and the bottom half representing the optimization phase, where orange represents the input text, blue represents the gradient, red represents the prompts
Appendix D Additional Results
D.1 Cost Analysis
@@
In this section, we evaluate the overhead of BATprompt in generating prompts. The primary overhead arises from the evaluation and generation processes during adversarial training. The total cost is represented by the following relation: $N\times(A+O)$, where $N$ is the number of iterations, $A$ represents the cost of the adversarial attack phase, and $O$ represents the cost of the optimization phase. When calling the LLM API, the cost is primarily determined by the number of tokens processed, including both input and output tokens. To estimate the cost of our method, we calculate the number of tokens required for executing tasks on three different datasets (XSum, Asset, and SST-5) across three types of tasks. This provides an understanding of the computational overhead associated with our approach. The results are in  Table 8
@@
Table 8: The cost of BATprompt in three dataset under two kinds of perturbations
From the results in the table, it can be observed that for relatively simple tasks (such as language understanding), the token consumption of the BATprompt algorithm is only 0.0258M. In contrast, for more complex tasks (such as text summarization), the token consumption remains at a low 0.0258M. After completing one round of BATprompt, the total token consumption is just 1.04M. This demonstrates that, once trained, BATprompt does not incur a large number of tokens. In comparison, Evoprompt consumes at least 4.20M tokens when converging on the SST-5 dataset, which has relatively simple input, significantly exceeding the token consumption of our method on XSum. This indicates that BATprompt offers substantial cost savings.
D.2 Language Understanding Task
Figure 12: In language understanding, specific indicators of 6 datasets on different 7 kinds of perturbations
In this section, we provide additional details of the data used in the experimental evaluation for the language understanding tasks. Results are in  Figure 13. We present the performance indicators of the prompts generated by BATprompt, alongside those obtained using other methods, across seven different perturbations on six datasets. The results show that BATprompt outperforms the other methods in most perturbation scenarios. Notably, in multi-label classification tasks such as SST-5 and TREC, BATprompt demonstrates stronger robustness to perturbations. This may be because binary classification tasks, with fewer labels and simpler task structures, are less affected by perturbations. In contrast, multi-label classification tasks are more complex, with a larger number of target categories, making them more vulnerable to perturbations. As a result, the prompts generated by BATprompt perform better in these more challenging scenarios.
D.3 Model Transferbility
Effectiveness on GPT-4o-mini. To demonstrate the universality of our method in different LLMs, we select the text summarization task and use GPT-4o-mini as the LLM backbone. The prompts generated by BATprompt, along with those generated by several other methods, are evaluated under different perturbations using Rouge-1, Rouge-2, and Rouge-L metrics. The results obtained are averaged, shown in the  Table 9. According to the experimental results, on different models, the prompts generated by BATprompt are still robust to different types of perturbations, and perform significantly better than other methods. In addition, we also tested the text simplification task and language understanding task on GPT-4o-mini. For the langugae-understanding task, We selected binary classification problem (CR) and multi-classification problem (SST-5) in sentiment classification datasets: as well as topic classification datasets. The specific experimental results are shown in the  Table 14 and  Table 12. In summary, our method also performs well on GPT-4o-mini.
@@ D.3 Model Transferbility
We plot the specific Rouge scores for each perturbation in the figure. The results show that, across all perturbations, our method demonstrates full robustness, outperforming the other prompt generation methods by a significant margin. Table 12 and Table 14 present the specific values of the prompt generated by BATprompt for the text simplification and language understanding tasks, respectively. In the text simplification task, BATprompt demonstrates optimal performance under various perturbations, with notable improvements in task scores, especially on GPT-4o-mini. Similarly, in the language understanding task, we selected several representative datasets, and the results show that our method retains robustness to perturbations across multiple datasets.
Figure 13: In the text summarization task, using GPT-4o-mini as backbone, the prompts produced by BATprompt and Rouge-1($\uparrow$), Rouge-2($\uparrow$), Rouge-L($\uparrow$) of the remaining methods on different disturbances.
Figure 14: In the text summarization task, using Llama2-7b as backbone, the prompts produced by BATprompt and Rouge-1, Rouge-2, Rouge-L of the remaining methods on different disturbances.
Effectiveness on Llama2-7b. We also transferred the instructions generated by BATprompt to a white-box model to assess the effectiveness of our prompts. For the text summarization task, we used Llama2-7b to compute Rouge-1, Rouge-2, and Rouge-L scores. To evaluate the robustness of our adversarial training prompts under various perturbations, we present the results in the Table 10. Our method still has optimal results. Further details on the text simplification and language understanding tasks can be found in the  subsection D.3.
@@
There are also additional details on the experiments conducted using Llama2-7b as the backbone for the text summarization task, as well as the experimental data for text simplification and language understanding. The results for the text summarization task are presented in the  Figure 14, where it is evident that BATprompt consistently outperforms baseline methods across all data items, as measured by Rouge-1, Rouge-2, and Rouge-L scores.
@@
BATprompt 41.60 40.54 40.13 42.82 41.83 43.54
Table 11: SARI($\uparrow$) values of the instruction generated by BATprompt under two types of perturbations, P1 and P2, under different perturbations on Asset in Llama2-7b
The SARI scores for the text simplification task on the Asset dataset are presented in the table. Similarly, the results demonstrate that the prompt generated by BATprompt outperform other methods across multiple types of perturbations.
Our experimental results on the language understanding task on the model of Llama2-7b are shown in the  Table 13. We calculated the average scores of each dataset under each perturbation. It can be seen that on the white-box model, the overall indicators of all datasets are lower. However, under the guidance of the instruction generated by BATprompt, the results of Llama2-7b have optimal values under both P1 and P2 perturbations. Through the above experiments, it can be seen that our method has sufficient robustness on both white-box and black-box.
@@
BATprompt 47.47 45.57 46.51 48.03 48.22 50.14
Table 12: SARI($\uparrow$) values of the instruction generated by BATprompt under two types of perturbations, P1 and P2, under different perturbations on Asset in GPT-4o-mini
@@
BATprompt 34.1 36.7 39.0 36.0 23.9 78.2 41.3
Table 13: Average score($\uparrow$) of the prompts from different method on six language understanding datasets using Llama2-7b. The table in the upper half is a perturbation of type P1, and the table in the lower half is a perturbation of type P2
@@
Table 14: The average of the scores($\uparrow$) of the instructions generated by BATprompt under two types of perturbations, P1 and P2, under different perturbations on the three datasets in GPT-4o-mini.
D.4 Perturbation Transferbility
In this experiment, we test prompts trained on a P1 perturbation against a P2 perturbation, prompts trained on a P2 perturbation against a P1 perturbation. The results in the Table 15. In the results, E-BATprompt refers to a prompt trained with a different type of perturbation. When tested on the current category, we observe that, except for the Rouge-1 and Rouge-2 scores of E-BATprompt in P1, which are higher than those of BATprompt, BATprompt achieves the best performance across the remaining metrics. Moreover, BATprompt still outperforms other methods significantly on unfamiliar perturbations.
Type Method Rouge-1 Rouge-2 Rouge-L
P1 EvoPrompt 17.52 3.14 14.97
E-BATprompt 21.35 4.59 16.10
BATprompt 21.25 4.44 16.66
P2 EvoPrompt 17.19 2.93 14.69
E-BATprompt 21.26 4.31 16.33
BATprompt 21.90 4.73 16.81
Table 15: The average of Rouge-1($\uparrow$), Rouge-2($\uparrow$) and Rouge-L($\uparrow$) of prompts trained by different perturbation on the text summarization task on GPT-3.5-turbo.
Appendix E Discussion
E.1 Number of iterations

### v3

@@ 6 Conclusion
In this paper, we highlight the vulnerability of existing automatic prompt word generation methods to input perturbations. To systematically evaluate this limitation, we introduce PertBench, a benchmark dataset that spans multiple tasks and incorporates a diverse set of perturbation types at the character, word, and sentence levels. PertBench enables a comprehensive assessment of prompt robustness under a wide range of input disturbances. Furthermore, we propose a novel prompt optimization method, PGO, that treats perturbation types as pseudo-gradient signals to guide the generation of more robust prompts in a gradient-free manner. Extensive experiments across various tasks and perturbation settings demonstrate the effectiveness and superior robustness of our approach compared to existing methods.
@@ 7 Limitation
We employed an efficient, training-free pseudo-gradient method. In the future, additional gradient-free techniques could be incorporated into our framework to further reduce generation costs and enhance performance. Moreover, although our method significantly outperforms existing approaches, its reliance on black-box models may limit the explainability of its superiority. Therefore, more advanced explainable AI (XAI) techniques may be required to better interpret this phenomenon.
Appendix A Details of Benchmark
Dataset Task Number
SST-2 {positive, negative} 1200
CR {positive, negative} 1200
MR {positive, negative} 1200
SST-5 {terrible, bad, okay, good, great} 1200
AG’s News {World, Sports, Business, Tech} 1200
TREC {Description, Entity, Expression, Human, Location, Number} 1200
ASSET Text Simplification 987
XSUM Text Summarization 720
Table 7: Statistics of our generating datasets for classification task and language generation task used in this work.
The Table. 7 shows the statistics of datasets we made for classification, text simplification, and text summarization tasks under different perturbations. Each dataset contains multiple sub-datasets, the total amount of generated data includes only perturbations to which the model is not robust for the given task. For example, the XSum dataset contains 7 sub-datasets corresponding to C2, C1, S1, W3, S3, S2, and W1 perturbations. For perturbations where the model is already robust, we constructed only 200 test samples per task. The robustness of the perturbation category is in the previous section.
A.0.1 Similarity between PertBench and Raw Data
@@
Table 8: The averagesemantic similarity and Levenshtein distance before and after attack in 3 tasks, 8 datasets.
In this part, we measure the similarity score between the generated benchmark and the original data set to prove that our dataset retains a high similarity with the original dataset. From the Table. 8, we can observe that for long text inputs, such as those in the XSum dataset, the Levenshtein distance and semantic similarity between the original and perturbed text remain above 98% for both types of perturbations (P1 and P2). In contrast, for shorter text inputs, such as those in the Asset dataset and the six classification tasks, even small changes can significantly impact the similarity measures. However, our experiments show that the Levenshtein distance similarity for P1 perturbed data remains above 90%, while the semantic similarity for P2 perturbed data stays above 80%, except for the SST-5. This indicates that our attack effectively preserves the essential information of the original sentence while introducing controlled perturbations.
Appendix B Experiments Setting
B.1 Hpyper Parameters
Our PGO algorithm is based on GPT-3.5-turbo for generation, with a total of 5 optimization iterations. During the perturbation phase, we set the number of iterative attacks to 3. In the combined perturbations and prompt optimization stages, we configured GPT-3.5-turbo with a Top-p value of 0.95 and a temperature of 1 to ensure both the robustness of the perturbations and the diversity of the generated outputs. For the testing phase, we set Top-p to 1 and temperature to 0, ensuring that the model produces consistent, fixed outputs.
B.2 Template
In this section, we give a complete template for the perturbation adding phase. Figure. 6, the prompt optimization phase. Figure. 7 provides a detailed illustration of the gradients generated during the optimization phase. Figure. 8 offers a comprehensive explanation of the prompt generation process, where gradients guide the LLMs, and highlights how prompt richness is enhanced through the rewriting process. The details of the specific implementation in LLMs are shown in Figure. 9.
@@ B.2 Template
Figure 9: PGO is based on the details of the LLMs implementation, with the left half representing the pseudo gradient perturbation phase and the right half representing the pseudo gradient optimization phase, where orange represents the input text, blue represents the gradient, red represents the prompts
Appendix C Additional Results
C.1 Experimental results for more methods
We compare our method against additional baselines (APE, InstructZero and INSTINCT) using GPT-4o-mini, and the results on both text generation and text understanding tasks are presented in the Table. 9, Table. 11 and Table. 12. It can be seen that our method has strong robustness on perturbed data and performs well in different tasks.
From the experimental results, we observe that the three baseline methods are not robust to perturbed text, with their performance consistently degrading after perturbations are introduced. Notably, the performance drop is more pronounced on multiple choice tasks such as TREC and SST-5. We believe this is because these methods do not explicitly provide an initial prompt, but instead attempt to infer it from the input and output. In the context of multiple choice tasks, this inference-based approach may struggle to effectively capture all options and accurately identify the target category, which likely contributes to the lower performance.
C.2 The stability of the results
To evaluate the performance stability of all the generated prompts, we selected the top five prompts and calculated the variance of their corresponding metrics, as well as the ratio of the variance to the mean. The results are presented in Table. 10.
From the results, we observe that for both summarization and simplification tasks, the standard deviation of ROUGE-1, ROUGE-2, ROUGE-L, and SARI scores for our generated prompts is consistently below 1.0, with most tasks exhibiting values under 0.5. Furthermore, the standard deviation-to-average ratio remains under 10%, indicating the stability of our prompts in evaluating text similarity. For classification tasks, we evaluated three datasets including: CR, SST-5 selected from the six datasets used in the paper. These datasets encompass binary classification, multi-class classification, sentiment classification, and topic classification. Our results show that the standard deviation-to-mean ratio remains below 10%, further demonstrating the robustness and stability of our prompt across diverse classification tasks.
APE 17.52 3.94 14.15 17.67 3.70 14.44 17.42 3.69 14.09 18.46 4.42 15.14
InstructZero 17.47 4.62 14.27 16.45 3.69 13.84 17.46 4.20 14.99 17.12 3.56 14.03
INSTINCT 17.01 4.34 14.33 15.44 3.40 12.86 15.63 3.18 12.96 17.61 3.83 14.17
PGO 18.79 4.38 15.05 18.20 4.06 14.92 17.78 3.98 14.71 18.96 4.33 15.45
Table 9: The Rouge-1($\uparrow$) score, Rouge-2($\uparrow$) score and Rouge-L($\uparrow$) score obtained by the prompts generated by the other methods for the task where the text summarization task is weak under the P2 class perturbation.
Type Rouge-1 Rouge-2 Rouge-L SARI CR SST-5
R1-std std/avg R2-std std/avg RL-std std/avg SARI-std std/avg CR-std std/avg SST5-std std/avg
P1 1.04 5.23% 0.23 6.37% 0.38 2.47% 0.40 0.8% 1.20 1.47% 1.63 3.55%
P2 0.64 3.30% 0.20 5.45% 0.32 2.05% 0.43 0.92% 3.23 3.79% 2.51 5.46%
Table 10: The average standard and ratio of standard and average.
C.3 Anomaly analysis
In Table. 4, we can see the results in Evoprompt on TREC dataset is near 0. This is an anomaly, we can analyze the reasons behind Evoprompt’s poor performance through specific examples on the TREC dataset. Our evaluation follows Evoprompt’s original prompt for the TREC dataset: Identity the inputs (explanations, entities, or humans) and use the outputs (numbers, descriptions, or entities) to answer the questions to make it easy understand for non-native English speakers. The task classifies questions into [’Description’, ’Entity’, ’Expression’, ’Human’, ’Location’, ’Number’]. However, Evoprompt’s prompt introduces a misleading causal relationship between input and output and lacks sufficient labels. Evoprompt uses Alpaca-7B for prompt generation and evaluation in classification tasks, excluding GPT-3.5-turbo. Our tests on other white-box models like Llama2-7B showed stable performance Table. 21. However, we analyze small-scale white-box models struggle with true semantic understanding, relying more on keyword extraction for classification. In contrast, GPT-3.5-turbo and GPT-4o-mini are more sensitive to prompt nuances, making them prone to misinterpretation if prompts are poorly optimized. Evoprompt’s prompts may mislead these models, causing errors, highlighting the need for transferable prompts—a strength of our approach.
C.4 Adding Suffix
In addition to demonstrating the effectiveness of our method, we also append suffixes directly after the initial prompts. These suffixes explicitly describe the type of perturbation present in the test data. We evaluate this approach across datasets from the three tasks on GPT-4o-mini, and the results are shown in Table. 14, Table. 13, Table. 15.
We can see that while intuitively, adding a suffix directly after the prompts seems like a reasonable approach, its effectiveness is limited. Our approach comprehensively outperforms adding a statement to initial prompts. Except for the C1 perturbation of the disturbed SST-5 dataset, where adding a statement slightly outperforms our method. We believe that this is because only adding suffixed may not be able to avoid the impact of perturbation, and the combination of original input and perturbation to explore the semantic space can obtain better results.
APE 45.36 40.96 43.83 44.42 43.78 44.43
InstructZero 33.16 32.95 30.12 25.24 20.16 27.14
INSTINCT 24.72 22.56 24.44 25.36 18.63 21.25
PGO 47.47 45.57 46.51 48.03 48.22 50.14
Table 11: The SARI($\uparrow$) score obtained by the prompts generated by the APE,InstructZero, INSTINCT for the task where the text simplification task.
Type Methods CR SST-2 MR SST-5 TREC AG’news Avg.
P1 APE 67.5 11.0 61.5 7.0 0.0 50.0 45.0
InstructZero 81.0 79.5 78.5 20.0 1.0 15.5 45.9
INSTINCT 58.5 35.0 33.5 2.0 0.0 31.5 26.8
PGO 86.8 75 84.8 44.5 65.3 82.0 73.0
P2 APE 73.4 13.2 62.0 12.6 0.6 48.0 35.0
InstructZero 74.2 82.8 76.2 18.6 0.6 15.2 44.6
INSTINCT 54.4 31.8 34.6 2.0 0.4 24.0 24.5
PGO 83.5 83.7 81.8 42.5 59.1 83.5 72.4
Table 12: Average score($\uparrow$) of the prompts from different method(APE, InstructZero, INSTINCT) on six classification datasets on GPT-4o-mini. The table in the upper half is a perturbation of type P1, and the table in the lower half is a perturbation of type P2.
Methods S1 W3 W1 S2 S3 W1
Add-suff 42.56 32.60 38.15 42.38 45.25 48.11
PGO 47.47 45.47 46.51 48.03 48.22 50.14
Table 13: SARI($\uparrow$) values of the instruction generated by PGO under two types of perturbations, P1 and P2, compared with adding suffix.
Type Method Rouge-1 Rouge-2 Rouge-L
P1 Add-suff 16.63 3.91 13.84
PGO 18.40 4.15 14.81
P2 Add-suff 16.45 3.76 13.73
PGO 18.43 4.19 15.03
Table 14: The average of Rouge-1($\uparrow$), Rouge-2($\uparrow$) and Rouge-L($\uparrow$) of prompts generated by different methods on the text summarization task
Method CR SST-5 TREC
P1.Pert P2.Pert P1.Pert P2.Pert P1.Pert P2.Pert
Add-suff 83.5 75.6 44.5 40.4 52.5 55.8
PGO 86.8 83.5 44.5 42.5 65.3 61.1
Table 15: The average of the scores($\uparrow$) of the instructions generated by PGO under two types of perturbations, P1 and P2, under different perturbations on the three datasets compared with adding suffix.
C.5 Cost Analysis
@@
In this section, we evaluate the overhead of PGO in generating prompts. The primary overhead arises from the evaluation and generation processes during Optimization. The total cost is represented by the following relation: $N\times(A+O)$, where $N$ is the number of iterations, $A$ represents the cost of the perturbation phase, and $O$ represents the cost of the optimization phase. When calling the LLMs API, the cost is primarily determined by the number of tokens processed, including both input and output tokens. To estimate the cost of our method, we calculate the number of tokens required for executing tasks on three different datasets (XSum, Asset, and SST-5) across three types of tasks. This provides an understanding of the computational overhead associated with our approach. The results are in Table. 16
@@
Table 16: The cost of PGO in three dataset under two kinds of perturbations
From the results in the table, it can be observed that for relatively simple tasks (such as classification), the token consumption of the PGO algorithm is only 0.0258M. In contrast, for more complex tasks (such as text summarization), the token consumption remains at a low 0.0258M. After completing one round of PGO, the total token consumption is just 1.04M. This demonstrates that, once trained, PGO does not incur a large number of tokens. In comparison, Evoprompt consumes at least 4.20M tokens when converging on the SST-5 dataset, which has relatively simple input, significantly exceeding the token consumption of our method on XSum. This indicates that PGO offers substantial cost savings.
Figure 10: In classification, specific indicators of 6 datasets on different 7 kinds of perturbations
C.6 Model Transferbility
Effectiveness on GPT-4o-mini. To demonstrate the universality of our method in different LLMs on the baseline, we select the text summarization task and use GPT-4o-mini as the LLMs backbone. The prompts generated by PGO, along with those generated by several other methods, are evaluated under different perturbations using Rouge-1, Rouge-2, and Rouge-L metrics. The results obtained are averaged, shown in the Table. 17. According to the experimental results, on different models, the prompts generated by PGO are still robust to different types of perturbations, and perform significantly better than other methods. In addition, we also tested the text simplification task and classification task on GPT-4o-mini. For the langugae-understanding task, We selected binary classification problem (CR) and multi-classification problem (SST-5) in sentiment classification datasets: as well as topic classification datasets. The specific experimental results are shown in the Table. 22 and Table. 20. In summary, our method also performs well on GPT-4o-mini.
@@ C.6 Model Transferbility
We plot the specific Rouge scores for each perturbation in the figure. The results show that, across all perturbations, our method demonstrates full robustness, outperforming the other prompt generation methods by a significant margin. Table. 20 and Table. 22 present the specific values of the prompt generated by PGO for the text simplification and classification tasks, respectively. In the text simplification task, PGO demonstrates optimal performance under various perturbations, with notable improvements in task scores, especially on GPT-4o-mini. Similarly, in the classification task, we selected several representative datasets, and the results show that our method retains robustness to perturbations across multiple datasets.
Figure 11: In the text summarization task, using Llama2-7b as backbone, the prompts produced by PGO and Rouge-1, Rouge-2, Rouge-L of the remaining methods on different disturbances.
Effectiveness on Llama2-7b. We also transferred the instructions generated by PGO to a white-box model to assess the effectiveness of our prompts. For the text summarization task, we used Llama2-7b to compute Rouge-1, Rouge-2, and Rouge-L scores. To evaluate the robustness of our optimized prompts under various perturbations, we present the results in the Table 18. Our method still has optimal results.
@@
There are also additional details on the experiments conducted using Llama2-7b as the backbone for the text summarization task, as well as the experimental data for text simplification and classification. The results for the text summarization task are presented in the Figure. 11, where it is evident that PGO consistently outperforms baseline methods across all data items, as measured by Rouge-1, Rouge-2, and Rouge-L scores.
@@
PGO 41.60 40.54 40.13 42.82 41.83 43.54
Table 19: SARI($\uparrow$) values of the instruction generated by PGO under two types of perturbations, P1 and P2, under different perturbations on Asset in Llama2-7b
The SARI scores for the text simplification task on the Asset dataset are presented in the table. Similarly, the results demonstrate that the prompt generated by PGO outperform other methods across multiple types of perturbations.
Our experimental results on the classification task on the model of Llama2-7b are shown in the Table. 21. We calculated the average scores of each dataset under each perturbation. It can be seen that on the white-box model, the overall indicators of all datasets are lower. However, under the guidance of the instruction generated by PGO, the results of Llama2-7b have optimal values under both P1 and P2 perturbations. Through the above experiments, it can be seen that our method has sufficient robustness on both white-box and black-box.
@@
PGO 47.47 45.57 46.51 48.03 48.22 50.14
Table 20: SARI($\uparrow$) values of the instruction generated by PGO under two types of perturbations, P1 and P2, under different perturbations on Asset in GPT-4o-mini
@@
PGO 34.1 36.7 39.0 36.0 23.9 78.2 41.3
Table 21: Average score($\uparrow$) of the prompts from different method on six classification datasets using Llama2-7b. The table in the upper half is a perturbation of type P1, and the table in the lower half is a perturbation of type P2
@@
Table 22: The average of the scores($\uparrow$) of the instructions generated by PGO under two types of perturbations, P1 and P2, under different perturbations on the three datasets in GPT-4o-mini.
Appendix D Discussion
D.1 Number of iterations